# Limits on the Axion–Photon Coupling from Chandrayaan-2

### v1

@@ V DATA ANALYSIS
Figure 4: 95% confidence level upper limits on $g_{a\gamma\gamma}$ from the three background subtracted solar x-ray spectrum as shown in Fig. 3. Also shown are the existing upper limits from the CAST experiment [22], the NuStar analysis of solar x-ray [26] and the projected upper limits from IAXO [40, 23] and ALPS-II [41] helioscope experiments.
VI RESULTS AND DISCUSSION
@@
With the analysis described above, the observation of soft solar x-ray from quiet-Sun by the XSM onboard Chandrayaan-2 yields the following 95% credible upper limit on the axion–photon coupling:
$g_{a\gamma\gamma}\lesssim(0.47\text{--}2.2)\times 10^{-10}\,\mathrm{GeV}^{-1},$ (10)
@@
for axion masses $m_{a}\lesssim 5\times 10^{-4}\,\mathrm{eV}$. Here the numerical spread reflects the variation due to the three background subtraction schemes adopted in this analysis: Case 1 (green solid line) gives $g_{a\gamma\gamma}\lesssim 2.2\times 10^{-10}\,\mathrm{GeV}^{-1}$ and is the most conservative limit, Case 2 (blue solid line) with realistic background subtraction gives $g_{a\gamma\gamma}\lesssim 1.2\times 10^{-10}\,\mathrm{GeV}^{-1}$, and Case 3 (red solid line) yields the most stringent value $g_{a\gamma\gamma}\lesssim 4.7\times 10^{-11}\,\mathrm{GeV}^{-1}$ and is the target/optimistic limit from our analysis. All these limits are shown in Fig. 4. The limits are effectively mass-independent across this low-mass range; sensitivity degrades for larger $m_{a}$ as coherence in the axion–photon conversion is lost when the axion mass exceeds the plasma frequency at the relevant solar-atmosphere heights.
@@
To facilitate comparison with ongoing and future laboratory efforts, we also overlay the limits from existing and future experiments. Shown in Fig. 4 are the existing limits from the CAST [22] experiment (gray shaded region) and the projected limits from the ALPS-II [41] (black dashed line) and IAXO [40, 23] (black dot-dashed line). For completeness, we also include in Fig. 4 the recent NuSTAR solar x-ray bound [26] derived from quiet-Sun observations, which provides the most stringent astrophysical limit to date. Although the most optimistic XSM curve (Case 3, red) is still weaker than the recent NuSTAR result of $6.9\times 10^{-12}\,\mathrm{GeV}^{-1}$ by roughly an order of magnitude, the XSM constraints are nevertheless competitive with CAST and, crucially, are complementary: XSM observed the entire solar disk simultaneously, reducing any systematic uncertainties associated with partial-disk observations [^1] The small effective collecting area of the XSM results in the almost order of magnitude difference between our and the NuSTAR ($A_{\rm eff}\simeq 800\,\mathrm{cm}^{2}$) result..
@@
We have presented all three background treatments for two reasons. Firstly, showing Case 1 – Case 3 makes explicit the degree to which the final sensitivity depends on background modeling; the spread quantifies the systematic uncertainty associated with background subtraction. Secondly including Case 3 (the red/target curve) indicates the best sensitivity one may realistically expect from this dataset under optimistic background control. In practice, Case 2 should be regarded as our baseline (realistic) constraint, Case 1 as the conservative bound, and Case 3 as the target limit that motivates further work.
@@
Theoretical uncertainties in the solar axion flux calculation are at the few-percent level in the regime relevant to this study and do not drive the overall error budget. The consistency of the limits across the three background treatments—differences remaining within a $\mathcal{O}(1)$ factor—indicates that instrumental and modeling systematics are subdominant to statistical limitations, which are set primarily by the small XSM effective collecting area. The XSM result is also consistent with astrophysical requirements from the measurement of solar neutrino flux and from helioseismology: axion-induced energy loss must remain a small fraction of the solar luminosity and our constraints are well within that regime.
@@
Finally, this remains an ongoing experiment. We therefore present the comparison between the three limits both to document current robustness and to define a clear target for future analyses: continued XSM observations, improved background characterization, and coordinated solar monitoring (photospheric/coronal magnetic mapping) can all push sensitivity toward the Case 3 target. More broadly, a future solar x-ray observatory, particularly around the next solar minimum, with a substantially larger collecting area and improved spectral resolution, would be capable of order-of-magnitude improvements, probing parameter space complementary to planned helioscopes. The Chanrayaan-2 XSM results thus motivate the further development of dedicated space-based solar observatories to search for axions and other physics beyond the Standard Model.
ACKNOWLEDGMENTS
@@
TK would like to thank Heerak Banerjee and Elisa M. Todarello for helpful insights regarding the numerical caluclations and data analysis parts. Additionally TK would like to thank Sougata Ganguly and Rohan Pramanick for helpful discussions and comments. TK also acknowledges support in the form of Senior Research Fellowship from the Council of Scientific & Industrial Research (CSIR), Government of India.

### v2

@@ Abstract
Figure 4: Shown are the 95% confidence level upper limits on $g_{a\gamma\gamma}$ derived from the three background-subtracted solar X-ray spectra in Fig. 3. Top Panel: Constraints in the $m_{a}–\,g_{a\gamma\gamma}$ plane for a fixed $g_{ae}=1.3\times 10^{-13}$, Bottom Panel: Constraints in the $g_{aee}-g_{a\gamma\gamma}$ plane obtained for a fixed $m_{a}=10^{-6}\,\mathrm{eV}$. Existing bounds and projected sensitivities are also shown for comparison.
@@
Results and discussion. With the analysis described above, the observation of soft solar x-ray from quiet-Sun by the XSM onboard Chandrayaan-2 yields the following 95% credible upper limit on the axion–photon coupling:
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}g_{a\gamma\gamma}\lesssim(0.50\text{--}2.26)\times 10^{-10}\,\mathrm{GeV}^{-1}$}, (10)
@@
for axion masses $m_{a}\lesssim 5\times 10^{-4}\,\mathrm{eV}$. Here the numerical spread reflects the variation due to the three background subtraction schemes adopted in this analysis: Case 1 (green solid line) gives $g_{a\gamma\gamma}\lesssim 2.26\times 10^{-10}\,\mathrm{GeV}^{-1}$ and is the most conservative limit, Case 2 (blue solid line) with realistic background subtraction gives $g_{a\gamma\gamma}\lesssim 1.29\times 10^{-10}\,\mathrm{GeV}^{-1}$, and Case 3 (red solid line) yields the most stringent value $g_{a\gamma\gamma}\lesssim 5.0\times 10^{-11}\,\mathrm{GeV}^{-1}$ and is the target/optimistic limit from our analysis. These limits are shown in the top panel of Fig. 4. The limits are effectively mass-independent across this low-mass range; sensitivity degrades for larger $m_{a}$ as coherence in the axion–photon conversion is lost when the axion mass exceeds the plasma frequency at the relevant solar-atmosphere heights.
@@
To facilitate comparison with current and future laboratory searches, in Fig. 4 we overlay existing experimental limits from CAST [CAST:2017uph] (gray shaded region) and projected limits for ALPS-II [Bahre:2013ywa] (black dashed line) and IAXO [Armengaud:2014gea, IAXO:2019mpb] (black dot-dashed line). We also include the recent bound from NuSTAR [Ruz:2024gkl] (magenta line), which is currently the most stringent astrophysical constraint. While the most optimistic XSM limit (Case 3) remains about an order of magnitude weaker than the NuSTAR bound of $7.3\times 10^{-12},\mathrm{GeV}^{-1}$, the XSM constraints are competitive with CAST and complementary to NuSTAR: XSM observes the full solar disk simultaneously, unlike NuSTAR, thereby avoiding systematics associated with partial-disk observations [^1] The smaller effective collecting area of XSM largely accounts for the difference relative to NuSTAR ($A_{\rm eff}\simeq 800,\mathrm{cm}^{2}$)..
Complementary to the mass-dependent limits above, the bottom panel of Fig. 4 shows constraints in the $g_{aee}$$g_{a\gamma\gamma}$ plane, obtained by fixing $m_{a}=10^{-6}\,\mathrm{eV}$, representative of the low-mass regime. The upper limits on $g_{a\gamma\gamma}$ strengthen monotonically with increasing $g_{aee}$, reflecting enhanced axion production in the solar interior. Existing bounds from horizontal-branch stars (brown shaded region), the white-dwarf luminosity function (olive shaded region), and CAST (gray shaded region) are shown for comparison [Barth:2013sma, Dent:2020jhf] and within the allowed parameter space, the Chandrayaan-2 XSM results do not significantly improve existing constraints on $g_{aee}$.
@@
We have presented all three background treatments for two reasons. Firstly, showing Case 1 – Case 3 makes explicit the degree to which the final sensitivity depends on background modeling; the spread quantifies the systematic uncertainty associated with background subtraction. Secondly, including Case 3 (the red/target curve) indicates the best sensitivity one may realistically expect from this dataset limited only by counting statistics, with an ideal background subtraction. The ideal background subtraction requires a comprehensive understanding of the various factors contributing to the detector background, including the cosmic X-ray background (CXB), the charged particle background of cosmic and solar origin, and secondary emission resulting from their interaction with the spacecraft structure. Apart from this, XSM, being a wide FOV instrument, is also affected by the variable bright astrophysical X-ray sources as well as the variable occultation of the CXB and particle background within the FOV by the Moon due to the altitude variation of the spacecraft. While Case 2 covers most of these aspects, some details, such as secondary emission from spacecraft structures and the time variability of all other X-ray sources within the FOV, which are very difficult to model, are not fully accounted for, and these will be necessary to reach the ideal Case 3. In practice, Case 2 should be regarded as our baseline (realistic) constraint, Case 1 as the conservative bound, and Case 3 as the target limit that motivates further work on accurate background modeling.
@@
Theoretical uncertainties in the solar axion flux are at the few-percent level and do not dominate the error budget. The consistency of the limits across the three background treatments—agreeing within an $\mathcal{O}(1)$ factor—shows that instrumental and modeling systematics are subdominant to statistical limitations, set primarily by the small XSM effective collecting area. The resulting constraints are also consistent with astrophysical requirements from solar neutrino measurements and helioseismology, which require axion-induced energy loss to remain a small fraction of the solar luminosity.
@@
Finally, this remains an ongoing experiment. The comparison of the three limits demonstrates current robustness and defines clear targets for future improvements: increased exposure, improved background modeling, and coordinated solar monitoring can push sensitivity toward the Case 3 benchmark. Looking ahead, a next-generation solar x-ray observatory—especially near solar minimum—with a substantially larger collecting area and improved spectral resolution could achieve order-of-magnitude gains, probing parameter space complementary to planned helioscopes. The Chandrayaan-2 XSM results thus motivate dedicated space-based solar observatories for searches for axions and other physics beyond the Standard Model.
@@
Acknowledgements. TK would like to thank Heerak Banerjee and Elisa M. Todarello for helpful insights regarding the numerical caluclations and data analysis parts. Additionally TK would like to thank Sougata Ganguly and Rohan Pramanick for helpful discussions and comments. The authors acknowledge the use of data from the Solar X-ray Monitor (XSM) on board the Chandrayaan-2 mission of the Indian Space Research Organisation (ISRO), archived at the Indian Space Science Data Centre (ISSDC). XSM was developed by Physical Research Laboratory (PRL) with support from various ISRO centers. TK also acknowledges support in the form of Senior Research Fellowship from the Council of Scientific & Industrial Research (CSIR), Government of India.